A case-control study examining the influence of sexual behaviour on Hepatitis B was conducted in a hospital between 2005 and 2008. The study recruited 229 Cases with Hepatitis B and 229 age- and sex-matched Controls. Both the Cases and Controls completed a questionnaire that inquired about the number of sexual partners they had in the past 2 years. Their responses were used to create the following exposure groups: 0 sexual partners (N=87), 1 sexual partner (N=257), and 2 or more sexual partners (N=114). Of the 229 Cases, 15 were in the 0 sexual partners group, 134 were in the 1 sexual partner group, and 80 were in the 2 or more sexual partners group.

Match the following questions to the appropriate answer:

  1. The prevalence of Hepatitis B in the 458 study participants was ____
  2. The prevalence of Hepatitis B in those with 2 or more partners was ____
  3. The odds ratio for Hepatitis B in the group with 1 sexual partner by comparison to the group with 0 sexual partners was ____
  4. The odds ratio for Hepatitis B in the group with 2 or more sexual partners by comparison to the group with 0 sexual partners was ____

a. 70%      
b. 50%      
c. 11.3      
d. 5.23
Answer #1
partner case control total
0 15 72 87
1 134 123 257
2 80 34 114
total 229 229 458

the prevalence of Hepatitis B in the 458 study participants was =(229/458)*100 =50%

The prevalence of Hepatitis B in those with 2 or more partners was =(80/114)*100 =70 %

The odds ratio for Hepatitis B in the group with 1 sexual partner by comparison to the group with 0 sexual partners was =(134/123)/(15/72)=5.23

The odds ratio for Hepatitis B in the group with 2 or more sexual partners by comparison to the group with 0 sexual partners was =(80/34)/(15/72)=11.3